A quantity of ideal gas at 12.0ºC and a pressure of 108 kPa occupies a volume of 2.47 m3. (a) How many moles of the gas are present? (b) if the pressure is now raised to 316 kPa and the temperature is raised to 31.0ºC, how much volume will the gas now occupy? Assume there are no leaks.

Grade:upto college level

1 Answers

areeba
4 years ago
P1V1÷T1= P2V2÷T2 Put the values 108×2.47÷12=316×V2÷31266.76÷12=316×V2÷3122.8=316×V2÷3122.8×31÷316=V2V2= 2.23 metrecube